# Unifying Demonstration Selection and Compression for In-Context Learning

### v2

@@ 6 Conclusion
This paper proposes UniICL, a parameter-efficient ICL framework that unifies demonstration selection, demonstration compression, and final response generation via a frozen LLM. Experimental results show that the generated virtual tokens substitute the 12$\times$ longer actual demonstrations with minimal time expenditure, scaling up the number of demonstrations from 4 to 64.
@@ 7 Limitations
Our study, while proposing an efficient unified ICL framework for demonstration compression and selection, still has limitations. Firstly, UniICL is limited to the realm of naive ICL leaving other advanced LLM prompting methods, e.g. Retrieval Augment Generation (RAG) and Chain-of-Thought (CoT) unexplored. Limited to the hardware, we employ the underlying LLM at a scale of 7 billion parameters. Larger-scale LLMs are welcome to enrich our findings in future studies.
Appendix A In-Domain Evaluation
@@
Table 5: The in-domain results and ablation studies on XSUM and CICERO. 1k represents the extending 1k window limitation, while others have a limitation of 512.
We conduct the zero-shot in-domain generation evaluation on the entire test set of XSUM and CICERO in Table 5 by compressing the latter half to virtual tokens and keeping the former unmodified. UniICL significantly outperforms the baselines, indicating the compressed virtual tokens can provide the original truncated information by recovering the cut-off parts after supervised fine-tuning. Although extending the window to 1k, Vicuna and BlueLM still underperform UniICL, indicating that compressed virtual tokens filter noise information to some extent.
Appendix B Results on BlueLM
We extra conduct experiments on BlueLM Team (2023) to verify the generality of UniICL. We demonstrate the result of understanding tasks in Table 6, of the generative tasks in Table 7.
Model #-shots CoLA-dev SST-2-dev IMDb Acc. BlueLM 0-shot 71.6 81.2 48.8 1-shot 69.6 82.6 64.8 2-shot 70.0 87.0 65.6 5-shot 70.5 88.6 68.7 UniICL 1-shot 69.6 81.2 65.4 2-shot 68.7 82.6 67.0 5-shot 71.7 87.0 70.4 UniICL 1-shot 69.8 80.0 62.0 2-shot 70.1 80.8 67.0 5-shot 71.8 85.6 69.6 8-shot 72.3 87.4 69.4 UniICL + $L_{ctr}$ 1-shot 70.1 80 69.6 2-shot 70.3 87.2 70.6 5-shot 71.1 89.2 71.0 8-shot 72.5 90.4 76.8
@@ Appendix B Results on BlueLM
Method #-shots XSUM ARXIV R-1 R-2 R-L R-1 R-2 R-L BlueLM 0-shot 15.0 3.6 10.4 30.9 7.7 24.7 1-shot 19.1 4.8 12.1 23.0 3.6 19.0 UniICL 1-shot 24.0 6.9 18.0 31.4 7.7 25.2 2-shot 25.0 7.3 18.8 30.8 7.3 24.8 5-shot 25.0 7.3 18.8 31.9 7.8 26.0 UniICL 1-shot 25.2 7.4 18.9 31.6 7.9 25.4 2-shot 25.4 7.6 19.1 31.9 8.0 25.6 5-shot 26.5 7.9 20.3 32.1 8.0 25.5
@@
Appendix C Datasets & Metrics
C.1 Datasets
Dataset # words (96,512] (512,1024] (1024,1536] XSUM - 10,000 4,697 CICERO 10,000 - - SUPER-NI - 10,000 7,000 XSUM (Ctr) 5,000
Table 8: The composition training set of UniICL. (m,n] represents the range of the number of words in each instance. XSUM (Ctr) is used for the second phase training in Equation 7.
Dataset Task In-Domain # Test # Demonstrations MS MARCO-dev Passage Ranking 6,980 - XSUM Text Summarization 1,500 204,045/20 ARXIV Text Summarization 1,500 203,037/20 CoLA-dev Lingustic Acceptability 1,041 67,349/20 SST2-dev Sentiment Classification 872 8,551/20 IMDb Sentiment Classification 1,500 25,000/20
Table 9: The details of involved evaluation datasets. -dev represents employing development set due to their test sets are inaccessible. # Demonstrations represent the number of demonstrations to be selected in high/low-resource ICL settings.
@@ C.1 Datasets
We mix up two public datasets for compression and selection augmentation training, described in Table 8. Additionally, UniICL achieves outstanding performance on out-domain evaluation, involving text summarization Narayan et al. (2018), passage ranking Nguyen et al. (2016), sentiment classification Maas et al. (2011); Socher et al. (2013), and linguistic acceptability Warstadt et al. (2018), more details referring to Table 9. UniICL selects demonstrations from its training set in high-resource ICL, and we fixed the number of candidate demonstrations to 20 for low-resource ICL evaluation.
C.2 Evaluation Metrics
@@
ROUGE Lin (2004) is a widely adopted metric in many generative tasks that evaluate how similar the generated hypothesis is to the golden label. Therefore, ROUGE is used in our experiments to evaluate the quality responses generated conditioned on compressed virtual tokens, and we report the F-1 scores of ROUGE-1, ROUGE-2, and ROUGE-L (abbreviated R-1, R-2, R-L in the following), and we employed the files2rouge  [^8] https://github.com/pltrdy/files2rouge.
library in practice. Following the previous works, we report the accuracy of close-ended evaluation and MRR@10 for passage ranking.

### v3

@@ 6 Conclusion
This paper proposes UniICL, a parameter-efficient ICL framework that unifies demonstration selection, demonstration compression, and final response generation via a frozen LLM, an adapter, and a learnable embedding. Experimental results prove the advantages of UniICL in both efficiency and effectiveness. Due to 12$\times$ demonstration compression, UniICL scales up the number of demonstrations from 4 to 64 within a 24 GB VRAM allocation. Finally, to avoid repeated compression of the same demonstration, UniICL configures a Demonstration Bank (DB, which significantly boosts model efficiency.
@@ 7 Limitations
Our study, while proposing an efficient unified ICL framework for demonstration compression and selection, still has limitations. Firstly, UniICL is limited to the realm of unmodified ICL, leaving other advanced LLM prompting methods, e.g., Retrieval Augment Generation (RAG) and Chain-of-Thought (CoT), unexplored. Limited to the hardware, we deploy the underlying LLM at a scale of 7 billion parameters. Larger-scale LLMs are welcome to enrich our findings in future studies.
8 Acknowledgement
I would like to express my sincere gratitude to all the authors and reviewers for their valuable contributions to this research. This work was supported by the National Natural Science Foundation of China (NSFC 62106165) and the Project Funded by the Priority Academic Program Development of Jiangsu Higher Education Institutions, China.
Appendix A Comparison with Existing Compression Methods
Methods Additional Compression # Trainable Train Compressor Tool Parameters Size LLMLingua Jiang et al. (2023) YES Pruning 7B 57k AutoCompressor Wingate et al. (2022) NO Soft Prompt 7B UNKNOWN ICAE Ge et al. (2023) YES Soft Prompt 70M 240k UniICL NO Soft Prompt 17M 47k
Table 8: Comparison among recent compression methods and UniICL. Compression Tool represents the involved compression technique of different methods. Train Size represents the size of the training datasets.
We present a comparison of training costs between UniICL and other recent compression methods in Tab. 8.
Appendix B In-Domain Evaluation
@@
Table 9: The in-domain results and ablation studies on XSum and CICERO. 1k represents the extended 1k window limitation, while others have a limitation of 512.
We conduct the zero-shot in-domain generation evaluation on the entire test set of XSum and CICERO in Tab. 9 by compressing the latter half to virtual tokens and keeping the former unmodified. UniICL significantly outperforms the baselines, indicating that the compressed virtual tokens can provide the original truncated information by recovering the cut-off parts after supervised fine-tuning. Although extending the window to 1k, Vicuna and BlueLM still underperform UniICL, indicating that compressed virtual tokens filter noise information to some extent.
Additionally, to quantify the performance gains brought by the learnable projection layer. We tune Vicuna and BlueLM with comparable parameters (17M) with LoRA, setting the rank to 32 in Tab. 9. UniICL still outperforms LoRA-adapted LLMs with a 512 window limitation, indicating that the truncation indeed brings performance degradation.
Appendix C Results on BlueLM
We also conduct experiments on BlueLM Team (2023) to verify the generality of UniICL. We demonstrate the result of understanding tasks in Tab. 10, of the generative tasks in Tab. 11.
Model #-shots CoLA-dev SST-2-dev IMDb Acc. BlueLM 0-shot 71.6 81.2 48.8 1-shot 69.6 82.6 64.8 2-shot 70.0 87.0 65.6 5-shot 70.5 88.6 68.7 UniICL 1-shot 69.6 81.2 65.4 2-shot 70.1 82.6 67.0 5-shot 71.8 87.0 70.4 UniICL 1-shot 70.1 80.0 62.0 2-shot 70.3 80.8 67.0 5-shot 71.1 85.6 69.6 8-shot 71.6 87.4 69.4 UniICL + $L_{ctr}$ 1-shot 68.9 80.0 69.6 2-shot 70.6 87.2 70.6 5-shot 71.5 89.2 71.0 8-shot 72.4 90.4 76.8
@@ Appendix C Results on BlueLM
Method #-shots XSum Arxiv R-1 R-2 R-L R-1 R-2 R-L BlueLM 0-shot 15.0 3.6 10.4 30.9 7.7 24.7 1-shot 19.1 4.8 12.1 23.0 3.6 19.0 UniICL 1-shot 24.0 6.9 18.0 31.4 7.7 25.2 2-shot 25.0 7.3 18.8 30.8 7.3 24.8 5-shot 25.3 7.4 19.1 31.9 7.8 26.0 UniICL 1-shot 24.7 7.2 18.5 31.0 7.5 24.9 2-shot 25.1 7.4 19.0 31.2 7.7 25.1 5-shot 26.3 7.6 20.0 31.5 7.9 25.3 UniICL + $L_{ctr}$ 1-shot 25.2 7.4 18.9 31.6 7.9 25.4 2-shot 25.4 7.6 19.1 31.9 8.0 25.6 5-shot 26.5 7.9 20.3 32.1 8.0 25.5
@@
Appendix D Supplementary Ablation on Llama2
AutoCompressor Wingate et al. (2022) and ICAE Ge et al. (2023) are built on Llama2-7B Touvron et al. (2023), which are soft-prompt-based methods similar to UniICL. Therefore, we evaluate UniICL with Llama2 as the backbone. As shown in Tab 12 and Tab. 13, UniICL achieves substantial improvements compared with unmodified Llama2 and outperforms ICAE and AutoCompressor demonstrated in Tab. 3.
Model #-shots CoLA-dev SST-2-dev IMDb Acc. Llama2 0-shot 73.4 93.0 85.3 1-shot 74.8 94.0 85.5 2-shot 75.6 94.9 87.8 5-shot 84.3 97.2 92.7 UniICL 1-shot 74.9 94.1 94.6 2-shot 75.9 95.1 96.1 5-shot 85.4 95.7 96.5
Table 12: The ICL results of understanding tasks with the backbone of Llama2.
Method #-shots XSum Arxiv R-1 R-2 R-L R-1 R-2 R-L Llama2 0-shot 27.4 7.6 20.1 32.9 8.9 29.2 1-shot 27.7 7.9 20.3 30.1 8.0 28.4 UniICL 1-shot 27.8 8.0 20.5 33.3 9.2 29.7 2-shot 28.4 8.6 21.3 34.0 9.4 30.3 5-shot 29.3 9.1 22.0 34.5 9.7 30.8
Table 13: The ICL results of generative tasks with the backbone of LLama2.
Appendix E Compression Ratio Selection on Different Tasks
Figure 9: Winrate of different compression ratios on out-of-domain evaluation in 1-shot settings.
We illustrate suitable ratio selection across four out-of-domain datasets in Fig. 9. For tasks with relatively short inputs, such as CoLA and SST2, UniICL tends to perform better with a compression ratio set to 4. While in IMDb and Arxiv, which are longer, UniICL performs better with higher compression ratios. UniICL with a 12$\times$ compression ratio substantially outperforms other settings on four datasets.
Figure 10: Winrate of UniICL with a fixed number of Memory Tokens.
Additionally, we are curious about whether it is necessary to introduce more demonstrations with a higher compression ratio. In Fig. 10, we find that the performance of compressing 2 demonstrations with a 12×12$\times$12 × ratio is stable and outperforms other settings across 3/4 datasets. 6$\times$ compression ratio with 1 demonstration compressed performs worst in general. When compressing 4 demonstrations with a 24×24$\times$24 × ratio, its performance is comparable, and it slightly outperforms the 12×12$\times$12 × ratio in SST2.
Appendix F Fixed Compression Ratio Training
Figure 11: The relative performance on in-domain and out-of-domain datasets, with UniICL trained with a fixed ratio. Out-of-domain evaluation applies 1-shot settings.
To verify the effectiveness of the dynamic sampled compression ratio of UniICL, we train models with more extensive fixed compression ratios and perform out-of-domain evaluation with the same ratio in Fig. 11. Results indicate that fixed compression ratios work better than dynamic sampled ratios in in-domain evaluation, but underperform in out-of-domain evaluation. We attribute this to the fixed compression ratio makes models exhibit over-fitting during training, and demonstration compression degrades to Prefix Tuning.
Appendix G Visualization of Memory Tokens
Figure 12: Cosine similarity between Memory Tokens (vertical axis) and original embeddings (horizon axis).
Figure 13: Attention scores on Memory Tokens in the first step generation. The vertical axis describes the 32 LLM layer, and the horizon axis indicates the number of Memory Tokens across different compression ratios. Above each figure, %/MT represents the average proportion of the attention score occupied by memory tokens in each LLM layer.
To explore how Memory Tokens work within UniICL across different compression ratios, we visualize the cosine similarity between Memory Tokens and original embeddings in Fig. 12 and attention scores of the first generation step in Fig. 13.
Intuitively, the 4$\times$ compression ratio should retain more information due to more Memory Tokens. However, as shown in Fig. 12(a), the cosine similarity is relatively sparser than the 4$\times$ compression ratio illustrated in Fig. 12(b). This tendency is aligned with the first step attention scores in Fig. 13(a). According to merely 0.3% average attention occupied in a generation, we can conclude that more Memory Tokens fail to provide models with more information. We attribute this phenomenon to the given semantic information being distributed over all Memory Tokens as models attend to each Memory Token equally in Fig. 13(a). Fewer Memory Tokens are enough to concentrate this information, represented as relatively concentrated similarity distribution in Fig. 12(b) and higher attention scores in Fig. 13(b), both of which indicate denser information retained. When the compression ratio becomes higher, such as 16 or 32, Memory Tokens become fewer and therefore sparse information retrained as shown in Fig. 12(c), Fig. 12(c), Fig. 13(c), and Fig. 13(c). This also provides an explanation for the slow performance degradation with ratios varying from 4 to 12 and drops sharply at 16 in Fig. 7.
Appendix H Datasets & Metrics
Datasets
@@ Appendix H Datasets & Metrics
We mix up three public datasets for compression and selection augmentation training, described in Tab. 1. The training set includes an instruction dataset, SUPER-NI, which we used to make UniICL respond to various instructions. Notably, we don’t perform an in-domain evaluation on SUPER-NI as it only contains a training set. After training, we extensively evaluate UniICL on out-of-domain evaluation, involving text summarization Narayan et al. (2018), passage ranking Nguyen et al. (2016), sentiment classification Maas et al. (2011); Socher et al. (2013), linguistic acceptability Warstadt et al. (2018), and a popular reasoning benchmark Hendrycks et al. (2020), more details referring to Tab. 2. MS MARCO is popularly used in Information Retrieval (IR), we use this dataset to evaluate the ability of UniICL to capture document-level information. Specifically, MMLU (Massive Multitask Language Understanding) is a comprehensive evaluation benchmark including 57 subjects from STEM, Humanities, Social Sciences, and Other fields. We use this benchmark to evaluate the reasoning ability of UniICL. UniICL selects demonstrations from its training set in high-resource ICL, and we fixed the number of candidate demonstrations to 20 for low-resource ICL evaluation.
Evaluation Metrics
@@
ROUGE Lin (2004) is a widely adopted metric in many generative tasks that evaluates how similar the generated hypothesis is to the golden label. Therefore, ROUGE is used in our experiments to evaluate the quality responses generated conditioned on compressed virtual tokens, and we report the F-1 scores of ROUGE-1, ROUGE-2, and ROUGE-L (abbreviated R-1, R-2, R-L in the following), and we employed the files2rouge  [^4] https://github.com/pltrdy/files2rouge. library in practice. Following the previous works, we report the accuracy of close-ended evaluation and MRR@10 for passage ranking.